Events A and B are mutually exclusive. Suppose event A occurs with probability 0.76 and event B occurs with probability 0.2. a. Compute the probability that A occurs but B does not occur. b. Compute the probability that either A occurs without B occurring or A and B both occur (If necessary, consult a list of formulas.)
Answer #1

(a)

Given:
P(A) = 0.76 (1)

P(B) = 0.2 (2)

A and B are mutually exclusive

To find P(AB :

By Theorem:

P(A)=P(Aigcap B)+P(Aigcap B^{C}) (3)

Since A and B are mutually exclusive:

P(Aigcap B)=0 (4)

Substituting (4) in (3), we get:

P(A) P(AB (5)

Substituting (1), we get from (5):

PLA )2c) = 0.76 (6)

So,

Answer is:

0.76

(b)

To find:

P(Aigcap B^{C} OR Aigcap B): (7)

From (4):

P(Aigcap B)=0 (8)

Substituting (8) in (7), we get:

P(Aigcap B^{C} OR Aigcap B)=P(Aigcap B^{C}) (9)

Substituting (9) in (6), we get:

P(Aigcap B^{C} OR Aigcap B)=0.76

So,

Answer is:

0.76